My Question
A piece of string 25 cm long goes exactly once around the edge of a plate in the form of a quadrant. Calculate the area of the quadrant.

Answer by fractalier(6550)

A quadrant (of a circle) has a perimeter of 2r + (1/2)(pi)r
If that equals 25, we can solve for r and then plug that into the formula for the area of it...A = (1/4)(pi)r^2
so we have
2r + (1/2)(pi)r = 25
r = 25 / (2 + (1/2)(pi))
A = (1/4)(pi)r^2
A = (1/4)(pi)(25 / (2 + (1/2)(pi)))^2
you can work out the exact value if needed...